OpenStudy (anonymous):

find the second and third derivatives of the following functions- (i) ax^2 + bx + c (ii) 2x/1-x (iii) 6x^4 - 3x - 4 (iv) 1+x/1-x(x is not equal to 1)

OpenStudy (anonymous):

1) Second Derivative : 2a Third Derivative = 0 2)Second Derivative=4/(1-x)^3 Third Derivative=12/(1-x)^4 3)Second Derivative=72x^2 Third Derivative=144x 4)Second derivative=4/(1-x)^3 Third Derivative=12/(1-x)^4